Project

General

Profile

0792014 Minutes

Attending: Evan, Michael, Mat

FD:

DB 2-9, 11-12
DIBLOCK 13 should go in by the end of week.

July 4th Rate drop - DCM out of sync (30 empty spils/event)
July 7 - Sporadic empty event behavior through until
- No dropped spills
- Correlated to increase in data rate, adding diblock,
- Slice Track ratio and difference metrics look very strange....

ND: Nitin will change the web page to make P2 the primary for now
NDOS: Timing peak is wide (Flaky TDU), NDOS lost it's lock at some point. (Does not affect other detectors. )


Siva - Sent mail to production group.......... 600k files for cosmics and numi
- Siva will generate test sample in "Mallard Release"
- Ask Barnali if she can validate.


Watchdog - I'm leaving on the 18th.